OpenStudy (mendicant_bias):

I'm having trouble with evaluating another inverse laplace transform. Posted in a moment in the comments section below.

OpenStudy (mendicant_bias):

\[L^{-1} \left\{ \vphantom{} \frac{1}{s^2+3s} \right\}\]

OpenStudy (mendicant_bias):

I'm not sure what I should do from here, but I think I'm going to fiddle with the denominator for a moment before I say "I don't know" and ask for help again.

OpenStudy (mendicant_bias):

It's a rather ugly solution, but I'm thinking of...well, no, nevermind, I can't do that. Assume that the Heaviside/Unit Step Function cannot be used and you can't "shift" anything. How should this be approached, then?

OpenStudy (mendicant_bias):

@perl

OpenStudy (perl):

use partial fractions

OpenStudy (perl):

1 / ( s^2 + 3s) = 1 / ( s ( 3+s) ) , now find partial fraction form

OpenStudy (perl):

i would like to use the draw button, but its not working :/ so i have to type things horizontally

OpenStudy (perl):

1 / ( s^2 + 3s) = 1 / ( s ( 3+s) ) = A / S + B / ( s + 3)

OpenStudy (perl):

so solve A, B for 1 / ( s ( 3+s) ) = A / S + B / ( s + 3)

OpenStudy (mendicant_bias):

Awesome! I just needed that initial idea, thank you.

OpenStudy (perl):

after you find A, B, then you can use your Laplace table (known transforms)

OpenStudy (sidsiddhartha):

hey we can use another property of inverse laplace here \[L^{-1}\frac{ f(s)}{ s }=\int\limits_{0}^{t}F(t).dt\\so~we~Have\\f(s)=\frac{ 1 }{ s(s+3) }\\now, ~L^{-1}\frac{ 1 }{ s(s+3) }=\int\limits_{0}^{t}L^{-1}\frac{ 1 }{ s+3 }.dt\\=\int\limits_{0}^{t}e^{-3t}.dt\] this makes it easiler @perl @Mendicant_Bias

OpenStudy (mendicant_bias):

Yeah, wait a minute, when I use partial fraction decomposition, I get a contradiction; can it still be done with PFD?

OpenStudy (mendicant_bias):

\[(A+B)s + 3 = 1\]This relationship must be true for all values of s, set s equal to zero. \[1 \neq 3\]

OpenStudy (mendicant_bias):

And @sidsiddhartha , did you just use convolution, or what? I'm totally unaware of that 1/s property and can't seem to find it anywhere else.

OpenStudy (sidsiddhartha):

no i did'nt use convolution if u multiply a function in S-Domain with s then it means-diffrentiation and if u multiply with 1/S then it means integration

OpenStudy (sidsiddhartha):

take a small example

OpenStudy (sidsiddhartha):

\[L^{-1}\frac{ s }{ s^2+1 }=?\]

OpenStudy (sidsiddhartha):

so here we can see \[\frac{ 1 }{ s^2+1 }~is~multiplied~with~s\]

OpenStudy (sidsiddhartha):

so first take inverse laplace of it